# Query Rewriting via LLMs

### v3

@@ Abstract
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms. The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer. For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison.
@@ 1 Introduction
The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools. Query optimizers are expected to, in principle, automatically remove such wasteful bloat while constructing efficient execution plans. However, in practice, they are often led astray by complex representations, resulting in poor response times.
SELECT t.Key, sum(t.Rating) AS PostRating,
(SELECT sum(b0.Rating)
FROM (SELECT p0.PostId, p0.BlogId, p0.Content,
p0.CreatedDate, p0.Rating, p0.Title,
b1.BlogId AS BlogId0, b1.Rating AS Rating0,
b1.Url, p0.day AS Key
FROM Posts AS p0 INNER JOIN Blogs AS b1
ON p0.BlogId = b1.BlogId
WHERE b1.Rating > 5) AS t0
INNER JOIN Blogs AS b0 ON t0.BlogId = b0.BlogId
WHERE t.Key = t0.Key ) AS BlogRating
FROM (SELECT p.Rating, p.day AS Key
FROM Posts AS p INNER JOIN Blogs AS b
ON p.BlogId = b.BlogId
WHERE b.Rating > 5) AS t
GROUP BY t.Key;
@@
SELECT p.day AS Key, SUM(p.Rating) AS PostRating,
SUM(b.Rating) AS BlogRating
FROM Posts AS p INNER JOIN Blogs AS b
ON p.BlogId = b.BlogId
WHERE b.Rating > 5
GROUP BY p.day;
@@
As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs. When this seemingly complex query is given to the current PostgreSQL engine (v16) [2], its execution plan essentially mimics the hierarchical structure of the query. This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete. However, the query can be equivalently rewritten (assuming NOT NULL column constraints and key-joins) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools in the hope that they may recommend performant alternatives. For this context, a viable S2S query transformer should satisfy the following criteria: (1) A wide range of queries should be covered in its scope; (2) The transformed query should be semantically equivalent to the original; (3) The rewrite should ideally improve performance, but at least not cause regression; and (4) The transformation overheads must be practical for deployment.
@@
A range of innovative Rule-based (e.g. Learned Rewrite [48]) and Model-based (e.g. Gen-Rewrite [28]) approaches have been proposed for S2S transformations. A recent vision paper [49] also advocates the use of LLMs for such query rewriting. The state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting. However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e transforming the query structure itself). In this paper, we address these lacuna and amplify the effectiveness of query rewriting. In particular, our query space-based transformations provide greater scope for candidate rewrites since only intent, and not implementation, is expressed, and this multiplicity of rewrites in turn results in a broader coverage of the underlying plan space.
@@
The LITHE Rewriter
We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention. As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance. Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$. Note that having the DBA in the loop is a common practice in commercial query advisory systems [14].
Our design of LITHE is based on a calibrated investigation of the suite of techniques described below:
Basic Prompts. We begin with a suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs. Interestingly, we find that more information is not necessarily better wrt rewriting quality, and that the best prompt granularity is query-specific. Moreover, this basic prompt ensemble was found to itself deliver performance similar to the SOTA techniques.
Database-sensitive Prompts. To help the model adapt to different query patterns and structures, we next introduce rules in the prompts. Our rules are invoked directly in query space, providing the LLM with the latitude to generalize the rule usage to a wide range of queries. This is in contrast to the hardwired and narrow rule application mechanisms (e.g. Calcite [10] rules) typically used in existing rewrite systems, which operate in plan space.
In particular, we work with two classes of rules – 1. Redundancy Removal Rules: These rules eliminate repeated and redundant computations of the same output; and 2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
Token Probability Driven Rewrites. The above rewrite options are restricted to the standard prompt interface. We also leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence. Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process. To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation. MCTS has previously been used to improve ordering of query rewrite rules [48] in the plan space, but our usage is new – it is geared to guide the LLM in exploring rewrites in the query space.
Semantic Equivalence Labels. After generating a rewrite, it is imperative to check for semantic equivalence to the original query. Unfortunately, the general problem of proving query equivalence is NP-complete [7]. We therefore take a twofold approach to check for query equivalence. First, the queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database. While these checks do not suffer from false negatives, they may (rarely) incur false positives. These tests can thus quickly weed out obviously non-equivalent rewrites.
A final check is then performed using logic-based equivalence-proving tools that provably work for a restricted classes of queries. Specifically, LITHE uses the recently developed QED tool [40] and SQLSolver [17] to verify rewrite correctness for queries within their scope. Whereas, for out-of-scope queries, result equivalence on the original database itself is used to test the equivalence.
After carrying out the above exercise, LITHE provides a verification label of semantic equivalence – provable if verified by a logic-based tool, or statistical otherwise. Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
Regression Identification. Finally, it is well-known that in practice there are often material discrepancies between optimizer cost predictions and actual execution times [29]. So it is possible that a promising rewrite as per the optimizer may actually regress at run-time as compared to the original formulation. To help guard against such possibilities, we have implemented heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
Experimental Evaluation
Our first set of experiments to evaluate LITHE’s performance is carried out on the industry standard TPC-DS benchmark [13], hosted on the PostgreSQL platform with GPT-4o used as the LLM. The evaluation focuses on slow queries taking more than a threshold time to complete. We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [48], LLM-R2 [27], GenRewrite [28], as well as a baseline LLM prompt [28]). The primary metrics are (a) reductions in optimizer-estimated costs, (b) run-time speedups, and (c) rewriting overheads. For LLM-based techniques, the number of tokens used is also monitored since the financial charges for LLM usage are typically dependent on this number. To understand the independent utility of the various components of LITHE, a systematic ablation study is carried out.
Generalizability.
In our second stage of experiments, we evaluate generalizability of the above outcomes in a variety of new scenarios, including (a) commercial database engines, (b) unseen database schemas, and (c) alternative LLM platforms.
@@
Results. Our experiments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract costs. In particular, for TPC-DS, LITHE constructed “highly productive” ($>1.5x$ estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number. Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
@@
We also evaluated whether the above cost reductions translated into real execution speedups. Here, we find that LITHE is indeed often substantively faster at run-time as well. Specifically, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison. Finally, our regression identification techniques were successful in guarding against “brittle” rewrites that are promising cost-wise, but eventually poor at run-time.
Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise application performance.
@@
    Evaluates rewrite quality over a broad range of database environments, demonstrating substantial benefits over both SOTA and the native optimizer.
@@ 2 LITHE Overview
The LITHE architecture, illustrated in Figure 3, consists of the following five-stage pipeline:
1. Prompt-based rewriting. This component consists of two modules: LLM Prompting and Syntax Verification. The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite. The LLM output’s syntax is then verified with the database engine’s parser– if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively (within a threshold) until a valid SQL query is obtained.
@@ 2.1 Performance Framework
The overall benefit provided by a rewriting tool is quantified by the number of CPR obtained on the slow query workload. Additionally, we also measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites. Finally, to assess the actual run-time benefits, we evaluate TSGM, the geometric mean (GM) of the response-time speedups obtained by these rewrites.
@@ 2.2 Query Micro-benchmark
Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sec 345). The TSGM performance is subsequently discussed in Section 7.
@@ 3 Basic Prompts
Prompt 4: The sequence of instructions in Prompt 3 is split into sub-prompts, and provided to the LLM in an iterative manner instead of all at once. The idea is to break down the complex instructions given in Prompt 3 into digestible steps that help the LLM focus on individual tasks.
@@ 4 Database-Sensitive Prompts
As discussed above, basic prompting needs to be improved on two fronts: (1) Ensuring productive rewrites where feasible; and (2) Maximizing the impact of these productive rewrites. To address these issues, we incorporate database domain knowledge. Specifically, we design a one shot-based prompting template, augmented with a set of database-aware rewrite rules. The rules are based on common practices followed by DBAs that are widely applicable, and augmented with precise instructions and useful examples to help guide the LLM in the rewriting process.
As a proof of concept, we explore two categories of rewrites here: (a) Rules that eliminate redundancy in the input queries; and (b) Predicate selectivity-based rules that implicitly guide, via query space reformulations, the query optimizer towards efficient query execution plans. Of course, this basic set of rules can be expanded further, but as shown by our experiments, even this minimal set is capable of delivering substantive improvements over a broad set of database environments.
@@ 4.1 Redundancy Removal
The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are available in the Section E. Note that this prompt template allows for only a single rule to be present in the prompt. This was a conscious design choice because LLMs are often overwhelmed by excessive information given in monolithic form. Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
@@
[…]
A natural question here would be whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied with the output of one prompt provided as input to the next, and so on. This approach would benefit queries with multiple types of redundancies. However, it also introduces significant computational overheads due to the vast number of possible rule application orders, making it expensive to explore. So, for simplicity, we have chosen to process them individually rather than cumulatively.
@@ 4.2 Selectivity-based Guidance
We now turn our attention to rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects. For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values. Rule R5 basically encodes this argument as a rule in Table 2. Similarly rule R6, which pre-filters tables that are involved in self-joins and have low selectivity filter and/or join predicates, is also used to guide a rewrite. Note that a specific instruction to not create explicit joins had to be added in rule R6. This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
@@ 5 Token Probability Driven Rewrite
A key challenge with LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs [20]. In order to have a robust approach for such cases, we take inspiration from the code generation literature [46]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements. As mentioned earlier, MCTS has been previously used for ordering rule applications in query rewrites [48], whereas our goal is to guide the LLM in exploring query rewrites.
This approach models the problem of query rewriting as a decision tree denoting a Markov Decision Process (MDP) [33]. The root node of the tree corresponds to the initial prompt. An edge from a parent node to a child represents a possible token generated by the LLM and is associated with a value denoting the probability of generating this token given the path taken thus far. Here, each edge can be considered as an action of the MDP. A node is considered terminal if the incoming edge corresponds to the “;” token, signalling end of the textual query.
The state of a node $n$ is represented by the partial rewrite created by following the path from the root to $n$ – it is obtained by concatenating the tokens on this path. The root’s state is an empty rewrite, and each terminal node’s state is a complete rewritten query.
@@
1:$root$          # Start State
@@
3:$\theta$              # Probability threshold for node expansion
4:$iter_{max}$     # Maximum number of iterations
5:$Potential$, $visits$, $V$leftarrow empty Map
@@
7:    $visits[root]$leftarrow $visits[root]+1$
8:    $n_{cur}$leftarrow $root$
@@
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
@@
15:    while $expand$ and ‘;’ $\notin n_{cur}.state$ do
16:        $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
@@
19:               $n_{new}\leftarrow$ new $Node$ with State $n_{cur}.state\cdot token$
@@
26:    $v\leftarrow$ ComputePotential($query$)

### v4

@@ Abstract
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms. The results show significant performance improvements for slow queries, over both SOTA rewriters and the native optimizer. For instance, with TPC-DS on PostgreSQL, the GM of runtime speedups was a high 13.2 over the native optimizer, whereas SOTA only gave 4.9.
@@ 1 Introduction
The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies [51, 2, 6], especially when machine-generated by modeling software such as ORM tools (e.g., Entity Framework [37], Hibernate [27]). Query optimizers are expected to, in principle, automatically remove such wasteful bloat while constructing efficient execution plans. However, in practice, they are often led astray by complex representations, resulting in poor response times.
@@
As a case in point, consider the blog-processing query [6] shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs. When this seemingly complex query is given to the current PostgreSQL engine (v16) [3], its execution plan essentially mimics the hierarchical structure of the query. This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete. However, the query can be equivalently rewritten (assuming NOT NULL column constraints and key-joins) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools for recommending performant alternatives. A viable S2S query transformer should cover a wide range of queries and ensure the rewrite is: (1) semantically equivalent to the original; (2) ideally performance-beneficial, but at least not causing regression; and (3) incurring overheads that are practical for deployment.
@@
A range of innovative Rule-based (e.g. Learned Rewrite [57]) and Model-based (e.g. Gen-Rewrite [34]) approaches have been proposed for S2S transformations. These state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting. However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e. transforming the query structure itself). In this paper, we address these lacuna and substantively amplify the effectiveness of query rewriting.
The LITHE Rewriter
We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries. As illustrated in the architectural diagram of Figure 3, LITHE takes the user query $Q^{U}$ as input and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement of $Q^{T}$, in terms of optimizer estimated cost; (b) a verification label indicating the semantic equivalence mechanism (provable or statistical) between $Q^{T}$ and $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance. Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$. Note that having the DBA in the loop is a common practice in commercial query advisory systems [17].
To help the LLM adapt to different query patterns and structures, we introduce database-aware rules into the prompts, which make use of the rich metadata (e.g., logical schema, selectivities) available in database environments (Section 4). Our rules are invoked directly in query space, providing the LLM with the latitude to generalize the rule usage to a wide range of queries. As shown later in our experiments, this is a powerful mechanism for ensuring performant rewrites across database environments. Furthermore, we leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence. Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process. To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation (Section 5).
LITHE also incorporates a suite of logic-based and result equivalence-based semantic equivalence tests to verify the correctness of the rewrites (Section 6.2). Finally, given that there are often material discrepancies between optimizer cost predictions and actual execution times [35], LITHE includes heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
@@
Results. Our experiments on benchmark environments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract (optimizer-estimated) costs. In particular, for TPC-DS, LITHE constructed “highly productive” (>1.5x>1.5x estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number. Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
@@
We also evaluated whether the projected cost reductions translated into real execution speedups. Here, we find that LITHE is indeed often substantively faster at run-time as well. Specifically, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison. Finally, our regression identification techniques were successful in guarding against “brittle” rewrites that are promising cost-wise, but eventually poor at run-time.
@@
    Evaluates rewrite quality over a range of database environments, demonstrating substantial performance benefits.
@@ 2 LITHE Overview
The LITHE architecture (Figure 3) comprises a five-stage pipeline:
1. Prompt-based rewriting. This component consists of two modules: LLM Prompting and Syntax Verification. The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite. The LLM output’s syntax is then verified with the database engine’s parser – if invalid, the error is returned to the LLM via an updated prompt asking for a correction. This goes on iteratively (within a threshold) until a valid SQL query is obtained.
@@ 2.1 Performance Framework
The overall benefit of a rewriting tool is quantified by the number of CPR obtained on the slow query workload. Additionally, we measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites. Finally, to assess the actual run-time benefits, we evaluate TSGM, the geometric mean (GM) of the response-time speedups obtained by these rewrites. These speedups are measured relative to the original query on the native optimizer.
@@ 2.2 Query Micro-benchmark
Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sections 345). The TSGM performance is subsequently discussed in Section 7.
@@ 3 Basic Prompts
Prompt 4: The sequence of instructions in Prompt 3 is split into sub-prompts, and provided to the LLM in an iterative manner instead of all at once. This breaks down the complex instructions of Prompt 3 into digestible steps that help the LLM focus on individual tasks.
@@ 4 Database-Sensitive Prompts
As discussed above, basic prompting needs to be improved on two fronts: (1) Ensuring productive rewrites where feasible; and (2) Maximizing the impact of these productive rewrites. Our discussions with industry experts revealed that query inefficiencies are most commonly attributed to redundant operations within the queries themselves. In fact, contemporary (non-LLM) optimizers often try to reduce redundancies during the initial logical optimization phase, prior to physical plan enumeration (e.g., redundant join elimination [9]). Furthermore, our analysis of execution plans for slow queries revealed that the optimizer often fails to substitute filter-related operators, even when beneficial to do so.
To address these issues, we incorporate database domain knowledge into the prompts. Specifically, we design a one shot-based prompting template, augmented with a set of database-aware rewrite rules. As a proof of concept, we explore two categories of rewrites here: (a) Rules that eliminate redundancy in the input queries; and (b) Predicate selectivity-based rules that implicitly guide, via query space reformulations, the query optimizer towards efficient query execution plans. Of course, this basic set of rules can be expanded further, but as shown by our experiments, even this minimal set is capable of delivering substantive improvements over a broad set of database environments.
The templates used for these prompts (Figure 5) apply only a single rule. This was a conscious design choice because LLMs can be overwhelmed by excessive information given in monolithic form [46]. A related question is whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied, thereby benefiting queries with multiple types of redundancies. However, the combinatorially large number of rule permutation sequences makes the selection process appear infeasibly expensive. Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best improvement. Despite this restrictive policy, highly performant rewrites are obtained in our evaluation (Section 7).
@@ 4.1 Redundancy Removal
The template for such rule-based prompts is shown in Figure 5(a) and includes an example rewrite to demonstrate the rule application to the LLM. Our empirical observation is that the precise wording of the rule instructions is not significant. Instead, what matters are the example rewrites, which ensure the LLM applies the rules effectively, and does not go off on tangential trajectories. The specific examples used with our rules are listed in Section D.
@@
Basic Prompts $\bigcup$ {R1, $\ldots$, R4} 7 6.85
@@ 4.2 Selectivity-based Guidance
We now consider rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects. For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively. Here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values. This observation is encoded in Rule R5 of Table 2. Similarly rule R6, which pre-filters tables that are involved in self-joins and with low selectivity filter and/or join predicates. Note that a specific instruction to not create explicit joins is added in rule R6. This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
@@ 5 Token Probability Driven Rewrite
A key challenge with LLMs is “hallucinations” – responses that range from being mildly incorrect to completely made up. This is often due to the output tokens having low confidence, which “confuse” the LLM and generate suboptimal outputs [23]. To have a robust approach for such cases, we take inspiration from the code generation literature [55]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite and performance improvements. While MCTS has been previously used for ordering rule applications in query rewrites [57], our goal is to guide the LLM in exploring these rewrites.
This approach models the problem of query rewriting as a decision tree denoting a Markov Decision Process (MDP) [41]. The root node of the tree corresponds to the initial prompt. Figure 7 illustrates one such example tree. An edge from a parent node to a child represents a possible token generated by the LLM and is associated with a value denoting the probability of generating this token given the path taken thus far. This is illustrated by the text (token) and number (probability), on the left and right side respectively, of an edge in Figure 7. Here, each edge can be considered as an action of the MDP. A node is considered terminal if the incoming edge corresponds to the “;” token, signaling end of the textual query.
The state of a node $n$ is represented by the partial rewrite created by following the path from the root to $n$ – it is obtained by concatenating the tokens on this path. The root’s state is an empty rewrite, and each terminal node’s state is a complete rewritten query. For example, the state of Node 4 in the example decision tree shown in Figure 7 is the partial rewrite ‘‘select * from’’.
@@
1:rootroot    # Start State
@@
[…]
5:PotentialPotential, visitsvisits, $V$ $\leftarrow$ empty Map
@@
7:  $visits[root]$ $\leftarrow$ $visits[root]+1$
8:  $n_{cur}$ $\leftarrow$ rootroot
@@
[…]
12:   $visits[n_{cur}]$ $\leftarrow$ $visits[n_{cur}]+1$
@@
15:  while expandexpand and ‘;’ $\notin n_{cur}.state$ do
16:   $tokens_{next}$, $P_{next}$ $\leftarrow$ Model($n_{cur},k$)
@@
19:      $n_{new}\leftarrow$ new NodeNode with State $n_{cur}.state\cdot token$
@@
26:  $v\leftarrow$ ComputePotential(queryquery)